# MC: Multi-concept Guidance for Customized Multi-concept Generation

### v2

@@ 4.2 Customized Multi-concept Generation Results
Qualitative evaluation. Figure 4 shows the generated images of our method and the baselines. Note that the problem definitions of [25, 15] slightly differ from the definition in Sec. 3.1, as they either require jointly training the two concepts or require training to merge the single-concept customized models. In our setting, the single-concept models are separately trained and directly used in inference. The baselines sometimes omit one of the specified concepts, e.g. the white chair for <sofa> and <chair> and the person for <person> and <cat>. Our method demonstrates higher fidelity to the reference images even compared to Custom Diffusion [25] that requires jointly training the two concepts, or Mix-of-Show [15] that requires training to merge the two single-concept customized models. Cones 2 [30] shows relatively low fidelity to the reference images, considering that it requires the least trained parameters.
Quantitative evaluation. We generate 16 samples per prompt with the same set of random seeds, resulting in a total of 19392 images for each method. As shown in Tab. 1, our method outperforms the baselines in terms of subject fidelity. Though [30] achieves a relatively higher CLIP-T score, it shows rather low subject fidelity. As mentioned in Sec. 4.1, CLIP-T is the similarity between the customized images and uncustomized text prompt. The prompt cat may be more close to some random cat image rather than a customized cat photo. A relatively high CLIP-T score indicates that our method has higher subject fidelity while consistent with the input text prompt.
@@ 4.3 Compositional Generation Results
Quantitative evaluation. We generate 16 samples per prompt with the same set of random seeds, resulting in a total of 4416 images for each method. As shown in Tab. 2, our method outperforms the baselines.
@@ 4.4 User Study
| Custom Diffusion | 3.66% | 7.77% | Composable Diffusion | 1.19% |
@@
We perform a user study to evaluate our method. For customized multi-concept generation, we show users the generated images of each method along with two reference images of the concept. Then the users are asked to select the image that is most consistent with the prompt and the image that best represents the concepts. For compositional generation, we show users the generated images of each method and ask them to choose the one most consistent with the prompt. The questionnaire contains 40 questions for the customization task and 30 questions for compositional generation. We collect 28 responses in total. As shown in Tab. 3, our method outperforms the baselines.
@@
| CLIP-T | CLIP-I | DINO | Full. | Min. | T-T | |
| Ours | 0.771 | 0.714 | 0.433 | 0.356 | 0.266 | 0.833 |
| Ours w/o first model | 0.764 | 0.711 | 0.425 | 0.349 | 0.263 | 0.809 |
| Ours w/o prefix | 0.632 | 0.736 | 0.476 | 0.354 | 0.266 | 0.822 |
| Ours w/o | 0.773 | 0.710 | 0.432 | 0.355 | 0.264 | 0.830 |
| Ours w/o , | 0.769 | 0.706 | 0.417 | 0.353 | 0.263 | 0.827 |
4.5 Ablation Study
To save computational expense, the ablation for customized multi-concept generation is carried out on a subset of CustomConcept101, which contains ten randomly selected compositions. Each composition is accompanied by 12 prompts. For compositional generation, we use the full benchmark dataset mentioned in Sec. 4.1.
The loss objectives. Without the two proposed loss terms, subject fidelity and prompt fidelity drop. Though harms prompt fidelity slightly, we include it to achieve a overall better result. As shown in Fig. 6, the fidelity to the reference images decreases when omitting the loss terms.
Architecture. As shown in Fig. 2, the first diffusion model is not equipped with any customized modules and is conditioned on prompt which describes the whole scene. We omit the first diffusion model then all the metrics decrease, as in the second row of Tab. 4. As shown in the third column of Fig. 6, mountains in the background are not generated without the first model.
Construction of sub-prompts. As mentioned in implementation details in Fig. 5, we append to as prefix to improve the consistency with the original input prompt, inspired by [17]. Without the prefix appended, the prompt fidelity drops, indicating the effect of this component to increase the consistency with the input prompt. The subject fidelity increases slightly as the model has more freedom to depict the concepts but with less control from the original prompt. In the fourth column of Fig. 6, mountains and city are not generated without the prefix.
@@ 5 Conclusion
In conclusion, we have proposed MC which allows combination of various heterogeneous single-concept customized models, enabling a more flexible customized multi-concept generation with improved fidelity. Multiple customized models are seamlessly integrated with Multi-concept Guidance to synthesize a natural composition of the multiple customized concepts, without additional training. MC can further be extended to elevate the compositional capabilities of text-to-image generation.
Appendix
@@
In Appendix 0.A, we show additional experimental results. In Appendix 0.B, we show more implementation details. In Appendix 0.C, Appendix 0.D and Appendix 0.E, we discuss the possible limitations, future work and societal impact of our work.
Appendix 0.A Additional Results
Figure 7 shows more qualitative comparisons of the proposed method and the baselines [30, 25, 15] on customized multi-concept generation. The prompts may contain other subjects besides the customized concepts, e.g. mountains and sunglasses. The fourth row shows the results on the composition of two concepts with similar appearance. The last three rows show results on editing the attribute of the customized concept, unreal composition of the concepts, and editing the style, respectively. Our method demonstrates a more satisfying effect.
Figure 8 shows qualitative comparisons of the customized multi-concept generation methods on composition of three concepts. Our method demonstrates more higher fidelity to the reference images.
Appendix 0.B Implementation Details
Table 5 shows the computational demands associated with each customized multi-concept generation method. Our measurements are conducted on a desktop equipped with i5-13600K and RTX 3090. The first column shows the storage requirement for a single concept. The comparatively lower storage demand of Mix-of-Show is partly due to its DropLoRA training strategy. This approach encourages the token embedding rather than the LoRA module to learn more of the concept, thus limiting its performance. We train all the models with fp16 mixed precision for 1000 steps, employing a batch size set of 2. The second column shows the training time of Custom Diffusion (Joint) on a composition of two concepts, and training time of other methods for a single concept. Custom Diffusion (Optimization) and Mix-of-Show require additional time for merging the single-concept models. Our method requires relatively longer inference time as it performs inference time optimization.
@@ Appendix 0.B Implementation Details
The coefficients are set to 0.8, 0.5, 0.4, respectively. For customized multi-concept generation, are set to 1.4, 5.6, 5.6, respectively. For compositional generation, are set to 5.6, 1.4, 1.4, respectively. The MCG is performed at the first 25 steps of the diffusion process. The Gaussian filter used to smooth the cross-attention maps has a kernel size of 3 and a standard deviation of 0.5.
| Method | Storage | Training Time | Additional Time | Memory Usage | Inference Time |
| Custom Diff. (Joint) [25] | 75MB | 9min | – | 4GB | 2s |
| Custom Diff. (Optim.) [25] | 75MB | 5min | 1min | 4GB | 2s |
| Mix-of-Show [15] | 5MB | 8min | 3min | 4GB | 2s |
| Cones 2 [30] | 5KB | 8min | – | 4GB | 2s |
| Ours | 38MB | 7min | – | 10GB | 8s |
Appendix 0.C Limitations
@@
A limitation of our method is that the architecture of multiple parallel diffusion models requires relatively larger memory usage, particularly when composing multiple customized concepts. This is partly due to our implementation. We literally maintain multiple instances of the same diffusion model in memory for the sake of simplicity. Addressing this limitation involves optimizing memory utilization by storing only a single instance of the diffusion model in memory, thereby enhancing memory efficiency.
@@
Despite MC enables users to compose multiple separately trained, even heterogeneous customized models, the customized models should be trained from the same diffusion model. Such limitation is inherited from [28].
Appendix 0.D Future Work
@@
In addition to addressing the limitations mentioned in Appendix 0.C, there exist several promising avenues for future research. [8, 48] delve into the realm of multi-concept customization for text-to-video generation. An intriguing prospect is to investigate the adaptability of our proposed MC to the domain of text-to-video generation. The existing benchmark dataset, CustomConcept101 [25], is currently limited to compositions of two concepts. To better evaluate the performance of methods in handling the composition of three or more concepts, it may prove beneficial to either develop a new dataset or expand upon the existing CustomConcept101.
@@
Appendix 0.E Societal Impact
First and foremost, MC empowers users to effortlessly generate visually captivating compositions reflecting their unique ideas and preference. Additionally, MC’s ability to enhance the capabilities of existing text-to-image diffusion models opens up new avenues for artistic exploration and innovation, potentially inspiring broader adoption and engagement in creative endeavors. However, MC may blur the lines between ethical and unethical image manipulation. Without proper guidance and ethical considerations, individuals may engage in harmful practices such as image forgery or digital impersonation. We advocate for the development of legal frameworks that address AI-generated content, including penalties for malicious use.

### v3

@@ 4.3 Customized Multi-concept Generation Results
Qualitative evaluation.
Figure 5 shows the generated images of our method and the baselines, containing compositions of two to four concepts. In the first row, we show the unreal composition of a cat wearing a jacket. Ours demonstrates a better consistency with both the cat and the jacket. In the fourth row, ours demonstrates superior editability compared to the baselines, successfully edits the chair to purple while not affecting the shape and color of the table. In the second to last row, we edit the artistic style of the image. In the last row, we show results of composing four concepts. The baselines failed to depict all the four concepts. The joint training of Custom Diffusion and model merging of Mix-of-Show sometimes make the concepts merge. As shown in the third and fourth row, the concepts would share some features. The Cones 2 and Textual Inversion have inferior concept consistency due to their architecture design that only an embedding is trained for each concept. For better consistency, more tunable parameters are needed. Ours avoids feature leakage between concepts in the training stage and the MCG provides a new way to compose them in the inference stage while reducing feature leakage.
Quantitative evaluation. We generate 16 samples per prompt with the same set of random seeds, resulting in a total of 33,024 images for each method. As shown in Tab. 1, our method outperforms the baselines in terms of prompt fidelity and subject fidelity. In the three and two concepts setting, Cones 2 demonstrates a higher prompt fidelity, which may come from its residual embedding design. The design reduces the impact of the customized concepts on the base model. However, a relatively small parameter size also hinders the learning of new concepts. Ours strikes a better balance between prompt fidelity and subject fidelity overall.
@@ 4.4 Compositional Generation Results
Quantitative evaluation. We generate 16 samples per prompt with the same set of random seeds, resulting in a total of 4,416 images for each method. As shown in Tab. 2, our method outperforms the baselines.
@@ 4.5 User Study
Custom Diff. | 3.66% | 7.77% | Composable Diff. | 1.19% |
@@
We perform a user study to evaluate our method. For customized multi-concept generation, we show users the generated images of each method along with two reference images of the concept. Then the users are asked to select the image that is most consistent with the prompt and the image that best represents the concepts. For compositional generation, we ask them to choose the one most consistent with the prompt. The questionnaire contains 40 questions for the customization task and 30 questions for compositional generation. We collect 28 responses in total. As shown in Tab. 3, our method outperforms the baselines.
4.6 Ablation Study
The ablation for customized multi-concept generation is carried out on a subset of MC++. For compositional generation, we use the full benchmark dataset mentioned in Sec. 4.2. We show ablation of our proposed MCG and attention grounding training in Tab. 4, Fig. 7.
@@ 4.6 Ablation Study
CLIP-T↑ | CLIP-I↑ | DINO↑ | Full.↑ | Min.↑ | T-T↑ | |
Ours | 0.789 | 0.716 | 0.434 | 0.356 | 0.266 | 0.833 |
0.784 | 0.712 | 0.426 | ||||
0.779 | 0.706 | 0.418 | 0.355 | 0.264 | 0.830 | |
0.440 | 0.539 | 0.356s | 0.353 | 0.263 | 0.827 |
4.7 Limitations
In Fig. 8, our method fails to compose two people, e.g. mixing their identities such as glasses. The attention maps of selected tokens are not sufficient for distinguishing the identities. With a refinement stage, the two people can be distinguished. Basically we give different weights to the regions of different concepts when merging the outputs of the unets. We do not include the refinement stage for the simplicity of our method. Details of the refinement stage are in suppl.
@@ 5 Conclusion
We have proposed MC2 which allows combination of various heterogeneous single-concept customized models, enabling a more flexible customized multi-concept generation with improved fidelity. Multiple customized models are seamlessly integrated with Multi-concept Guidance to synthesize a natural composition of the multiple customized concepts. MC2 can further be extended to elevate the compositional capabilities of text-to-image generation.
Supplementary Material
@@
In Sec. 6, we show the details of the MC++ benchmark. In Sec. 7, we show more implementation details, including hyperparameter setting and corresponding ablation. In Sec. 8, we discuss the refinement stage mentioned in Sec. 4.7 of the main paper and more possible limitations. In Sec. 9, we show additional results. In Sec. 10 and Sec. 11, we discuss future work and societal impact of our work.
6 MC++ Benchmark
We adapt CustomConcept101 [26] to compositions of three and four subjects. The original CustomConcept101 is a dataset of 101 concpets with 3-15 images for each concept. For evaluating multi-concept customization, the CustomConept101 contains prompts for 101 compositons of two subjects. To facilitate the evaluation of multi-concept customization on more than two subjects, we collect prompts for 57 compositions of three subjects, and 14 compositions of four subjects from the original CustomConcept101. Each composition has 12 prompts. Basicly we follow the procedure of [26] to get the prompts. We first used ChatGPT [7] to propose the prompts then manually filter them to get the final set of prompts.
7 Implementation Details
We choose LoRA as our single-concept customization model. We adopt a popular community implementation of LoRA [48]. LoRA modules are trained for linear layers and conv layers in text encoder and unet with rank set to 16. We do not finetune the token embeddings. All reference images are captioned as photo of a <concept name>. Each concept has a unique <concept name> defined by CustomConcept101, e.g. pet_dog1. All LoRAs are trained for 1000 steps with batch size set to 2, learning rate set to 1e-4. LoRA scale is set to 0.7 for merging the LoRA parameters into the diffusion model during inference.
@@ 7 Implementation Details
Here we list the hyperparameters used in our method. In Eq. (7), is set to 0.8. In Eq. (8), we schedule the learning rate linearly. It starts from 20 then decays to 10 linearly. We take a single gradient step per diffusion time step. In Eq. (9), are set to for customized multi-concept generation, set to for compositional generation. In Eq. (10), are set to 0.5, 0.4. The MCG is performed at the first 25 steps of the diffusion process. The Gaussian filter used to smooth the cross-attention maps has a kernel size of 3 and a standard deviation of 0.5. Figure 9 shows ablation of the hyperparameters. The controls the balance between and . The controls the balance between the uncustomized model and the customized models. The red box marks our recommended hyperparameters.
About the implementation of the baselines, we basically follow their official repo except Cones 2 [31]. It was implemented based on SD 2.1 rather than SD 1.5. As our experiments are based on SD 1.5, we adapt the official code to SD 1.5. However, we found their default hyperparameter sub-optimal for SD 1.5. As shown in Fig. 10, when the is set to default value 1e-3, the model failed to learn the <flower> concept. The is the coefficient of the regularization term for training a Cones 2 model. A large hinders the learning of the model. We set to 1e-5 in the experiments.
8 Limitations
As mentioned in Sec. 4.7 of the main paper, we add a refinement stage for the composition of multiple concepts that share similar features. When two concepts share similar features, it may be difficult to tell them apart based on the corresponding cross-attention maps. Here we propose an algorithm for extracting masks for the concepts based on the cross-attention maps. The goal is to propose a binary mask with soft boundary for each concept. Algorithm 1 shows how to extract masks for two subjects. We visualize the masks in Fig. 11. In 3 and 4, we rescale and smooth the attention maps. In 5, 6, 7 and 8, we detect the overlap of the two attention maps with a threshold. In 9, we binarize the attention maps to get the masks. From 10 to 15, we dilate the masks and assign the dilated area to the proper subject based on distance transform. Finally we perform Gaussian blur to get a soft boundary. The masks are then used in Eq. 14 for merging the outputs of multiple diffusion models, which is identical to [5].
Note that the masks are generated on the fly and we replace the Eq. (9) in the main paper with Eq. 14. As shown in the Fig. 11, with the refinement stage, the dog is correctly generated.
@@ 8 Limitations
Another limitation of our method is that the architecture of multiple parallel diffusion models requires relatively larger memory usage, particularly when composing multiple customized concepts. This is partly due to our implementation. We literally maintain multiple instances of the same diffusion model in memory for the sake of simplicity. Addressing this limitation involves optimizing memory utilization by storing only a single instance of the diffusion model in memory, thereby enhancing memory efficiency.
@@
Despite MC2 enables users to compose multiple separately trained, even heterogeneous customized models, the customized models should be trained from the same diffusion model. Such limitation is inherited from [29].
9 Additional Results
Figure 12 shows more qualitative comparisons of the proposed method and the baselines [31, 26, 16] on customized multi-concept generation. The baselines sometimes omit one of the specified concepts, e.g. the white chair for <sofa> and <chair> and the person for <person> and <cat>. Our method demonstrates higher fidelity to the reference images even compared to Custom Diffusion [26] that requires jointly training the two concepts, or Mix-of-Show [16] that requires training to merge the two single-concept customized models. Cones 2 [31] shows relatively low fidelity to the reference images, considering that it requires the least trained parameters. Our method demonstrates a more satisfying effect. Figure 13 shows qualitative comparisons of the customized multi-concept generation methods on composition of three concepts. Our method demonstrates more higher fidelity to the reference images.
Figure 14 shows more qualitative comparisons of the compositional generation methods. Our method demonstrates better consistency with the input text prompts compared to the baselines [8, 29, 28].
10 Future Work
@@
In addition to addressing the limitations mentioned in Section 8, there exist several promising avenues for future research. [9, 51] delve into the realm of multi-concept customization for text-to-video generation. An intriguing prospect is to investigate the adaptability of our proposed MC2 to the domain of text-to-video generation.
@@
11 Societal Impact
First and foremost, MC2 empowers users to effortlessly generate visually captivating compositions reflecting their unique ideas and preference. Additionally, MC2’s ability to enhance the capabilities of existing text-to-image diffusion models opens up new avenues for artistic exploration and innovation, potentially inspiring broader adoption and engagement in creative endeavors. However, MC2 may blur the lines between ethical and unethical image manipulation. Without proper guidance and ethical considerations, individuals may engage in harmful practices such as image forgery or digital impersonation. We advocate for the development of legal frameworks that address AI-generated content, including penalties for malicious use.